# Multi-Objective min-max Online Convex Optimization

### v1

@@ 4.2 Guarantee
The main ingredient of the proof as well as the inspiration for Algorithm 1 is the following decomposition (19) of the expected min-max static regret (18). Recall the definition of
$\Lambda_{t}(x,\theta):=\theta_{1}f_{t}(x)+\theta_{2}g_{t}(x).$
where $\theta=(\theta_{1},\theta_{2})\in\Delta_{2}$.
By adding and subtracting two terms, $\sum_{t=1}^{T}\Lambda_{t}(x_{t},\theta_{t})$ and $\min_{x\in\mathcal{X}}\sum_{t=1}^{T}\Lambda_{t}(x,\theta_{t})$ inside the expectation in the RHS of (18), we write the expected min-max static regret (18) as
${\bar{R}}_{T}(\mathcal{A})=\sup_{\mathcal{D}}\{{\mathbb{E}}\{R_{1}\}+{\mathbb{E}}\{R_{2}\}+{\mathbb{E}}\{R_{3}\}\}\leq\sup_{\mathcal{D}}\{{\mathbb{E}}\{R_{1}\}\}+\sup_{\mathcal{D}}\{{\mathbb{E}}\{R_{2}\}\}+\sup_{\mathcal{D}}\{{\mathbb{E}}\{R_{3}\}\},$ (19)
@@
$R_{3}=\min_{x\in\mathcal{X}}\sum_{t=1}^{T}\Lambda_{t}(x,\theta_{t})-C_{\mathsf{OPT}}=\min_{x\in\mathcal{X}}\sum_{t=1}^{T}\Lambda_{t}(x,\theta_{t})-\min_{x\in\mathcal{X}}\max_{\theta\in\Delta_{2}}\sum_{t=1}^{T}\Lambda_{t}(x,\theta).$
The interpretation of $R_{1}$ is that it is the regret for the gains version of $2$-Experts Problem Cesa-Bianchi and Lugosi (2006), where the two function sequences $f_{t},g_{t}$ are the gains of two experts, with respect to an optimal $\theta^{\star}\in\Delta_{2}$ for fixed actions $x_{t}$ across time slots. This motivates the use of Hedge algorithm Cesa-Bianchi and Lugosi (2006) (gains version) in our proposed algorithm that is tailor-made for updating $\theta_{t}$ for fixed actions $x_{t}$.
Similarly, $R_{2}$ is the static regret for an OCO minimization problem relative to static optimal action $x^{\star}$ for a single surrogate objective function sequence of $\Lambda_{t}(x,\theta_{t})=\theta_{t1}f_{t}(x)+\theta_{t2}g_{t}(x)$ for a fixed $\theta_{t}$. Hence using an OGD algorithm is natural for this purpose as done in the algorithm for each time $t$, with a fixed weight $\theta_{t}$.
@@
We will bound $R_{1}$ and $R_{2}$ sample path wise for each realization of $f_{t},g_{t}$, while for $R_{3}$ we will bound ${\mathbb{E}}\{R_{3}\}$. In particular, we will show the following that is proven in the Appendix.
@@
Because of the connection described above, the proof for $R_{1}$ in Lemma 5 and for $R_{2}$ in Lemma 6 follows standard analysis Hazan (2019) for Hedge and OGD, respectively. Bounding $R_{3}$ needs more care and we prove it in Appendix C where we critically exploit the i.i.d. input model.
Combining Lemma 5, 6 and 7, and the regret decomposition (19), we get that the expected min-max static regret (12) of Algorithm 1 with
$\eta_{\theta,t}\leq\sqrt{\frac{2\ln 2}{B^{2}t}},\qquad\ \text{and}\qquad\eta_{x,t}=\frac{D}{G\sqrt{t}}$ (20)
is
${\bar{R}}_{\mathcal{A}}(T)=O(\sqrt{T}),$ (21)
proving the result. ∎
Following remarks are in order.
@@
When there are $K$ loss function sequences $f_{t}^{k},k=1,\dots,K$ that are chosen i.i.d. from $\mathcal{D}$, define $\theta\in\Delta_{K}$, (where $\Delta_{K}$ is the $K$-dimensional simplex) and $\Lambda_{t}(x,\theta)=\sum_{k=1}^{K}\theta_{k}f_{t}^{k}$. Then the expected min-max static regret of $\mathcal{A}$ is
${\bar{R}}_{\mathcal{A}}(T)=\sup_{\mathcal{D}}{\mathbb{E}}\left\{\max_{\theta\in\Delta_{K}}\sum_{t=1}^{T}\Lambda_{t}(x_{t},\theta)-T\cdot\min_{x\in\mathcal{X}}\max_{\theta\in\Delta_{K}}\theta^{\top}\mu(x)\right\}$ (22)
where $\mu(x)={\mathbb{E}}\{f_{t}^{1}(x),\dots,f_{t}^{K}(x)\}$.
With this new definition of $\theta$ and $\Lambda_{t}$, Algorithm 1 remains as is with Hedge choosing $\theta_{t}\in\Delta_{K}$ while $x_{t}$ the single action chosen by OGD. Assuming all $f_{t}^{k},k=1,\dots,K$ satisfy Assumption 1 and 2, the only change in the guarantee with $K>2$ compared to Theorem 4 with $K=2$ is because of Lemma 5 that is now $O(\sqrt{T\log K})$ Hazan (2019) since there are $K$ experts corresponding to the $K$ loss function sequences $f_{t}^{k},k=1,\dots,K$ . The OGD guarantee of Lemma 6 stays the same since $\|\nabla\Lambda_{t}(x_{t},\theta)\|\leq G$ as before. Moreover, Lemma 7 holds as is. Thus, the overall expected min-max static regret of Algorithm 1 with $K$ loss function sequences is
$O(\sqrt{T\log K}).$ (23)
@@
We note that Algorithm 1 is well motivated even while considering the worst case/adversarial input. The technical problem in its analysis with the worst case/adversarial input is that Lemma 7 that is the critical result fails when considering the adversarial input. We highlight this difficulty in Appendix F and show that the ‘best possible’ upper bound on $R_{3}$ is $R_{3}=O(T^{3/2})$ with the worst case/adversarial input for Algorithm 1 by exploiting just the properties of the Hedge algorithm which is insufficient for achieving sub-linear regret.
@@
[Extension to Strongly Convex $f_{t}$’s and $g_{t}$’s] One limitation of our proposed algorithm and the proof is that because of the use of the Hedge algorithm for updating $\theta_{t}$, the regret $\sup_{\mathcal{D}}\{R_{1}\}$ can at best be $O(\sqrt{T})$. Thus, unlike OCO, where OGD algorithm achieves logarithmic regret in $T$ when $f_{t}$’s are strongly convex, our algorithm’s regret is only $O(\sqrt{T})$ even when both $f_{t},g_{t}$ are strongly convex. Is this a limitation of the algorithm or of that of the problem itself remains an interesting unresolved question.
5 Improving the result on GlobalExpertsProblem with the i.i.d. input Even-Dar et al. (2009)
Let $\mathcal{X}=\Delta_{K}$, the $K$-dimensional simplex, and writing $f_{t}^{k}(x)=x_{k}\ell_{k}(t)$, where $x_{k}$ is the $k$-th coordinate of $x\in\mathcal{X}$, $\ell_{k}(t)$ is the loss corresponding to the $k$-th expert, and $k=1,2\dots,K$, we get the GlobalExpertsProblem Even-Dar et al. (2009) with $K$ experts in the i.i.d. input case is a special case of the i.i.d. min-max OCO problem (18).
Thus, from (23), we get that the expected min-max static regret of Algorithm 1 is $O(\sqrt{T\log K})$ for the GlobalExpertsProblem with i.i.d. input. In comparison, Even-Dar et al. (2009) derived a min-max static regret (5) guarantee of $O(\sqrt{T}\log K)$ with adversarial input for the following algorithm called MULTI:
For $K=2$,
$x_{1}(t+1)=x_{1}(t)+\frac{x_{2}(t)\ell_{2}(t)-x_{1}(t)\ell_{1}(t)}{\sqrt{T}},$
and $x_{2}(t)=1-x_{1}(t)$. For general $K$, the $K=2$ algorithm was used recursively. Compared to our Algorithm 1, the MULTI algorithm Even-Dar et al. (2009) is very compute intensive since it recursively uses an algorithm for $K=2$ and requires the knowledge of time horizon $T$.
@@ 6 Extension to Bandit Feedback
We now analyze the expected min-max OCO problem (18) under the bandit feedback, where at time $t$, once the online algorithm $\mathcal{A}$ chooses its action $x_{t}$, it only observes the scalar values $\{f_{t-1}^{k}(.)\}_{k=1}^{K}$ at one (one-point model) or two points (two-point model) and does not receive any gradient information.
Algorithm 2 Bandit-MinMax (unified one-point/two-point variant)
1:Feasible Set $\mathcal{X}\subset\mathbb{R}^{d}$, smoothing parameters $\{\delta_{t}\}$, step-sizes $\{\eta_{x,t}\},\{\eta_{\theta,t}\}$, feedback type $\mathsf{mode}\in\{\text{one-point},\text{two-point}\}$
2:Initialize $w_{1}=\mathbf{1}_{K}$, $\theta_{1}=w_{1}/\|w_{1}\|_{1}$, choose $x_{0}\in X$
3:Define $\Lambda_{t}(x,\theta_{t})=\sum_{k}\theta_{t,k}f_{t}^{k}(x)$
4:for $t=1,\dots,T$ do
5:  Sample $u_{t}$ uniformly from unit sphere $\mathbb{S}^{d-1}$
6:  if $\mathsf{mode}=\text{one-point}$ then
7:   Play $\tilde{x}_{t}=\text{Proj}_{\mathcal{X}}(x_{t-1}+\delta_{t}u_{t})$ and observe $\{f_{t}^{k}(\tilde{x}_{t})\}_{k=1}^{K}$
8:   Gradient estimator:
$\widehat{\nabla}_{t}=\frac{d}{\delta_{t}}\Lambda_{t}(\tilde{x}_{t},\theta_{t})\,u_{t}$
9:  else if $\mathsf{mode}=\text{two-point}$ then
10:   Query $x_{t-1}+\delta_{t}u_{t}$ and $x_{t-1}-\delta_{t}u_{t}$, observe $\{f_{t}^{k}(x_{t-1}\pm\delta_{t}u_{t})\}_{k=1}^{K}$
11:   Gradient estimator:
$\widehat{\nabla}_{t}=\frac{d}{2\delta_{t}}\Big(\Lambda_{t}(x_{t-1}+\delta_{t}u_{t},\theta_{t})-\Lambda_{t}(x_{t-1}-\delta_{t}u_{t},\theta_{t})\Big)u_{t}$
12:  end if
13:  Gradient step: $y_{t}=x_{t-1}-\eta_{x,t}\widehat{\nabla}_{t}$,    $x_{t}=\text{Proj}_{\mathcal{X}}(y_{t})$
14:  Hedge update: $w_{t+1,k}=w_{t,k}\exp\big(\eta_{\theta,t}f_{t,k}(x_{t})\big)$,    $\theta_{t+1}=w_{t+1}/\|w_{t+1}\|_{1}$
15:end for
6.1 Algorithm
For this bandit model, we consider a simple adaption of Algorithm 1 by replacing its OGD part with the bandit information counterpart, while keeping the Hedge part as it is, given by Algorithm 2. Moreover, the analysis also follows the same script by replacing the guarantees by their respective bandit counterparts.
6.2 Guarantee
@@
Under Assumption 1 and 2,
@@
    (One-point feedback) With $\delta_{t}=t^{-1/4}$, $\eta_{x,t}=D/(Gd^{1/2}t^{3/4})$, and $\eta_{\theta,t}=\sqrt{\log K}/(B\sqrt{t})$, the expected min-max static regret satisfies
    ${\bar{R}}_{\mathcal{A}}(T)\leq c_{1}\,d^{1/2}GD\,T^{3/4}\;+\;c_{2}B\sqrt{T\log K},$
    for universal constants $c_{1},c_{2}>0$.
@@
    (Two-point feedback) With $\delta_{t}=t^{-1/2}$, $\eta_{x,t}=D/(Gd^{1/2}\sqrt{t})$, and $\eta_{\theta,t}=\sqrt{\log K}/(B\sqrt{t})$, the expected min-max static regret satisfies
    ${\bar{R}}_{\mathcal{A}}(T)\leq c_{3}\,d^{1/2}GD\,\sqrt{T}\;+\;c_{4}B\sqrt{T\log K},$
    for universal constants $c_{3},c_{4}>0$.
6.3 Proof of Theorem 8
Recall the decomposition (19)
Bounding $R_{1}$ and $R_{3}$.
For both one-point or two-point bandit feedback, the Hedge guarantee remains the same as in Lemma 5. In particular, with $\eta_{\theta,t}=\sqrt{\log K}/(B\sqrt{t})$, sample path wise,
$R_{1}\leq B\sqrt{2T\log K},$
and similar to Lemma 7,
${\mathbb{E}}\{R_{3}\}\leq 0.$
The only change is in bounding $R_{2}$.
Bounding $R_{2}$ (one-point feedback). The one-point estimator
$\widehat{\nabla}_{t}=\frac{d}{\delta_{t}}\Lambda_{t}(\tilde{x}_{t},\theta_{t})\,u_{t}$
is unbiased for the gradient of the $\delta_{t}$-smoothed loss, with variance at most $O(dB^{2}/\delta_{t}^{2})$. Using $\delta_{t}=t^{-1/4}$ and $\eta_{x,t}=D/(Gd^{1/2}t^{3/4})$, the bandit-OGD analysis Flaxman et al. (2005) yields
$R_{2}\leq c_{1}\,d^{1/2}GD\,T^{3/4},$
for some universal constant $c_{1}$.
Combining the three bounds we get
${\bar{R}}_{\mathcal{A}}(T)\leq c_{1}\,d^{1/2}GD\,T^{3/4}\;+\;2B\sqrt{2T\log K}.$
Bounding $R_{2}$ (two-point feedback). With two function evaluations per round, the estimator
$\widehat{\nabla}_{t}=\tfrac{d}{2\delta_{t}}\Big(\Lambda_{t}(x_{t-1}+\delta_{t}u_{t},\theta_{t})-\Lambda_{t}(x_{t-1}-\delta_{t}u_{t},\theta_{t})\Big)u_{t}$
is unbiased with variance $O(dB^{2}/\delta_{t}^{2})$. Choosing $\delta_{t}=t^{-1/2}$ and $\eta_{x,t}=D/(Gd^{1/2}\sqrt{t})$, the two-point bandit-OGD analysis Agarwal et al. (2010); Duchi et al. (2015); Shamir (2017) gives
$R_{2}\leq c_{3}\,d^{1/2}GD\,\sqrt{T},$
for some universal constant $c_{3}$. Thus, with two-point feedback,
${\bar{R}}_{\mathcal{A}}(T)\leq c_{3}\,d^{1/2}GD\,\sqrt{T}\;+\;2B\sqrt{2T\log K}.$

### v3

@@ 6.2 Guarantee
The main ingredient of the proof as well as the inspiration for Algorithm 1 is the following decomposition (10) of the expected min-max static regret (8). Let $\Delta_{K}$ be the $K$-dimensional simplex. Recall the definition of $\Lambda_{t}(x,\theta):=\theta_{1}f_{t}(x)+\theta_{2}g_{t}(x)$, where $\theta=(\theta_{1},\theta_{2})\in\Delta_{2}$. By adding and subtracting two terms, $\sum_{t=1}^{T}\Lambda_{t}(x_{t},\theta_{t})$ and $\min_{x\in\mathcal{X}}\sum_{t=1}^{T}\Lambda_{t}(x,\theta_{t})$ inside the expectation in the RHS of (8), we write the expected min-max static regret (8) as
${\bar{R}}_{\mathcal{A}}(T)=\sup_{\mathcal{D}}\{{\mathbb{E}}\{R_{1}\}+{\mathbb{E}}\{R_{2}\}+{\mathbb{E}}\{R_{3}\}\}\leq\sup_{\mathcal{D}}\{{\mathbb{E}}\{R_{1}\}\}+\sup_{\mathcal{D}}\{{\mathbb{E}}\{R_{2}\}\}+\sup_{\mathcal{D}}\{{\mathbb{E}}\{R_{3}\}\},$ (10)
@@
and
$R_{3}=\min_{x\in\mathcal{X}}\sum_{t=1}^{T}\Lambda_{t}(x,\theta_{t})-{\bar{C}}_{\mathsf{OPT}}=\min_{x\in\mathcal{X}}\sum_{t=1}^{T}\Lambda_{t}(x,\theta_{t})-T\cdot\min_{x\in\mathcal{X}}\max_{\theta\in\Delta_{2}}\theta^{\top}\mu(x).$
The interpretation of $R_{1}$ is that it is the regret for the gains version of $2$-Experts Problem Cesa-Bianchi and Lugosi (2006), where the two function sequences $f_{t},g_{t}$ are the gains of two experts, with respect to an optimal $\theta^{\star}\in\Delta_{2}$ for fixed actions $x_{t}$ across time slots. This motivates the use of Hedge algorithm (gains version) in our proposed algorithm that is tailor-made for updating $\theta_{t}$ for fixed actions $x_{t}$.
Similarly, $R_{2}$ is the static regret for the OCO problem relative to static optimal action $x^{\star}$ for a single surrogate objective function sequence of $\Lambda_{t}(x,\theta_{t})=\theta_{t1}f_{t}(x)+\theta_{t2}g_{t}(x)$ for a fixed $\theta_{t}$. Hence using an OGD algorithm is natural for this purpose as done in the algorithm for each time $t$, with a fixed weight $\theta_{t}$.
@@
We bound $R_{1}$ and $R_{2}$ sample-path wise for each realization of $f_{t},g_{t}$, while for $R_{3}$ we will bound ${\mathbb{E}}\{R_{3}\}$ as follows whose proofs are provided in Appendix E, F, and G, respectively
@@
Combining Lemma 6, 7 and 8, and the regret decomposition (10), we get that the expected min-max static regret ${\bar{R}}_{\mathcal{A}}(T)=O(\sqrt{T})$ (8) with $\eta_{\theta,t}\leq\sqrt{\frac{2\ln 2}{B^{2}t}},\text{and}\ \eta_{x,t}=\frac{D}{G\sqrt{t}}$. ∎
@@
For the $\alpha$-fair problem, Si Salem et al. (2023) derived a regret guarantee of $O(\alpha\sqrt{T})$, which becomes vacuous for our regime of interest since $\alpha$-fair problem is equivalent to our min-max setting when $\alpha\rightarrow\infty$. Compared to Si Salem et al. (2023), Algorithm 1, has an $\alpha$-independent regret guarantee. It is tempting to use Algorithm 1 for the $\alpha$-fair problem for finite $\alpha$, however, as we point out in Remark 11, the $\alpha$-fair problem with finite $\alpha$ and $\alpha\rightarrow\infty$ are fundamentally different.
@@
When there are $K$ loss function sequences $f_{t}^{k},k=1,\dots,K$ that are chosen i.i.d. from $\mathcal{D}$, define $\theta\in\Delta_{K}$, (where $\Delta_{K}$ is the $K$-dimensional simplex) and $\Lambda_{t}(x,\theta)=\sum_{k=1}^{K}\theta_{k}f_{t}^{k}$. With $\mu(x)={\mathbb{E}}\{f_{t}^{1}(x),\dots,f_{t}^{K}(x)\}$, analogous to (8), the expected min-max static regret of $\mathcal{A}$ is
${\bar{R}}_{\mathcal{A}}(T)=\sup_{\mathcal{D}}{\mathbb{E}}\left\{\max_{\theta\in\Delta_{K}}\sum_{t=1}^{T}\Lambda_{t}(x_{t},\theta)-T\cdot\min_{x\in\mathcal{X}}\max_{\theta\in\Delta_{K}}\theta^{\top}\mu(x)\right\}.$ (11)
With this new definition of $\theta$ and $\Lambda_{t}$, Algorithm 1 remains as is with Hedge choosing $\theta_{t}\in\Delta_{K}$ while $x_{t}$ the single action chosen by OGD with respect to the surrogate loss function $\Lambda_{t}(x,\theta_{t})$ with fixed $\theta_{t}$. Assuming all $f_{t}^{k},k=1,\dots,K$ satisfy Assumption 1 and 2, the only change in the guarantee with $K>2$ compared to Theorem 5 with $K=2$ is because of Lemma 6 that is now $O(\sqrt{T\log K})$ Hazan (2019) since there are $K$ experts corresponding to the $K$ loss function sequences $f_{t}^{k},k=1,\dots,K$ . The OGD guarantee of Lemma 7 stays the same since $\|\nabla\Lambda_{t}(x_{t},\theta)\|\leq G$ as before. Moreover, Lemma 8 holds as is. Thus, the overall expected min-max static regret of Algorithm 1 with $K$ loss function sequences is $O(\sqrt{T\log K}).$
@@
[Extension to Strongly Convex $f_{t}$’s and $g_{t}$’s] One limitation of our proposed algorithm and the proof is that because of the use of the Hedge algorithm for updating $\theta_{t}$, the regret $\sup_{\mathcal{D}}\{R_{1}\}$ can at best be $O(\sqrt{T})$ with stochastic i.i.d. input without any further assumptions. Thus, unlike OCO, where OGD algorithm achieves logarithmic regret in $T$ when $f_{t}$’s are strongly convex, our algorithm’s regret is only $O(\sqrt{T})$ even when both $f_{t},g_{t}$ are strongly convex. Is this a limitation of the algorithm or of that of the problem itself remains an interesting unresolved question.
@@ 7 Extension to Bandit Feedback
We now analyze the expected min-max OCO problem (8) under bandit feedback: at each time $t$, after choosing $x_{t}$, the algorithm observes only the scalar losses $\{f_{t-1}^{k}(\cdot)\}_{k=1}^{K}$ at one point (one-point model) or two points (two-point model), and receives no gradient information. For this bandit model, we consider a simple adaptation of Algorithm 1 by replacing its OGD part with the bandit information counterpart, while keeping the Hedge part as it is, given by Algorithm 2. Exploiting the modular nature of the proof in Section 6, the results follow almost immediately.
Algorithm 2 Bandit-min-max
1:Input:Feasible set $\mathcal{X}\subset\mathbb{R}^{d}$, parameter $\{\delta,\rho\}$, stepsizes $\{\eta_{x,t}\},\{\eta_{\theta,t}\}$, $\mathsf{mode}\in\{\text{one-point},\text{two-point}\}$ Initialize $w_{1}=(1,1,\dots,1)_{K}$, $\theta_{1}=w_{1}/\|w_{1}\|_{1}$, choose $x_{1}\in(1-\rho)\mathcal{X}$
2:for $t=1,\dots,T$ do Sample $u_{t}\sim\mathrm{Unif}(\mathbb{S}^{d-1})$
3:  if $\mathsf{mode}=\text{one-point}$ then Query $\tilde{x}_{t}=x_{t}+\delta u_{t}$  Observe $\{f_{t}^{k}(\tilde{x}_{t})\}_{k=1}^{K}$$\widehat{\nabla}_{t}=\frac{d}{\delta}\Lambda_{t}(\tilde{x}_{t},\theta_{t})u_{t}$ where $\Lambda_{t}(\tilde{x}_{t},\theta_{t})=\sum_{k=1}^{K}\theta_{t,k}f_{t}^{k}(\tilde{x}_{t})$$\ell_{t,k}=f_{t}^{k}(\tilde{x}_{t})$
4:  elseQuery $x_{t}^{+}=x_{t}+\delta u_{t}$, $x_{t}^{-}=x_{t}-\delta u_{t}$  Observe $\{f_{t}^{k}(x_{t}^{+})\}_{k=1}^{K}$ and $\{f_{t}^{k}(x_{t}^{-})\}_{k=1}^{K}$$\widehat{\nabla}_{t}=\frac{d}{2\delta}\big(\Lambda_{t}(x_{t}^{+},\theta_{t})-\Lambda_{t}(x_{t}^{-},\theta_{t})\big)u_{t}$$\ell_{t,k}=\tfrac{1}{2}\left(f_{t}^{k}(x_{t}^{+})+f_{t}^{k}(x_{t}^{-})\right)$
[…]
6:end for
@@
Let $\mathsf{r}$} be such that $r\mathcal{B}_{1}\subseteq\mathcal{X}\subseteq D\mathcal{B}_{1}$ and $\mathcal{B}_{1}$ is the unit ball.
@@
    (One-point feedback) With $\delta=T^{-1/4}$, $\eta_{x,t}=D/(Gd^{1/2}t^{3/4})$, $\rho=\delta/{\mathsf{r}}$, and $\eta_{\theta,t}=\sqrt{\log K}/(B\sqrt{t})$, expected min-max static regret ${\bar{R}}_{\mathcal{A}}(T)\leq c_{1}\,d^{1/2}GD\,T^{3/4}\;+\;c_{2}B\sqrt{T\log K}$, for constants $c_{1},c_{2}>0$.
@@
    (Two-point feedback) With $\delta=T^{-1/2}$, $\rho=\delta/{\mathsf{r}}$, $\eta_{x,t}=D/(Gd^{1/2}\sqrt{t})$, and $\eta_{\theta,t}=\sqrt{\log K}/(B\sqrt{t})$, expected min-max static regret satisfies ${\bar{R}}_{\mathcal{A}}(T)\leq c_{3}\,d^{1/2}GD\,\sqrt{T}\;+\;c_{4}B\sqrt{T\log K}$, for constants $c_{3},c_{4}>0$.
Proof of Theorem 9 can be found in Appendix I.